three +20uC charges are placed along a straight line,succesive charges being 2m apart.what is the force on the charge on the right end ?

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
force=(k*20*20/4^2 +k*20*20/2^2)*10^(-12)N
=(400*k/16 +400*k/4)*10^(-12)N
=125*k=125*9*10^9 *10^-12=1.125N


Sher Mohammad
b.tech iit delhi
faculty askiitians